OpenStudy (anonymous):

Find the difference ten and two-ninths minus seven and two-thirds. Be sure to reduce to simplest form.

OpenStudy (anonymous):

i will give medals

OpenStudy (anonymous):

and fan

OpenStudy (automaticloveletter):

2/9 - 23/3 (7 and 2/3 to an improper fraction) gcf= 3 2/2=1 9/2=3 1/3-23/3= -22/3 (I could be wrong)

OpenStudy (anonymous):

thank you for the help